OpenStudy (fluffybear):

Just a little confused? What is the value of the variable that makes the statement true? a + (–7) + 3 = 0 A. –4 B. –3 C. 4 D. 7

OpenStudy (fluffybear):

At first i thought the answe was 7 but then i saw the 3? ?_?

OpenStudy (anonymous):

Well first add the number 7 to each side

OpenStudy (anonymous):

Which would be what? -7+7=? 0+(-7)=?

OpenStudy (fluffybear):

well i think -7+7 would equal 0 and uhhh.. i really dont know lol

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Am I meant 0+7=?

OpenStudy (anonymous):

Sorry about the confusion

OpenStudy (anonymous):

a+3=7

OpenStudy (anonymous):

subtract 3 from both sides and you have your answer

OpenStudy (fluffybear):

thank you :)

OpenStudy (anonymous):

You're welcome. Have a nice day. :)

OpenStudy (triciaal):